How do you set up a capital budget projection on a fictional company using Excel? Estimating output, output prices, revenues. When will expenses be paid? Chose a discount rate and calculate the PV of future cash flows.
Answer #1

Capital budgeting may involves purchasing raw material, assets and arranging required services or anything which has monitory cost associated. Companies usually take projects which have positive margin over the cost.

In NPV approach we discount the after-tax cash flows by the weighted average cost of capital and if derived value is greater then the incurred cost then project considered to be profitable.

To set up a budget for any project, first we should calculate the associated implementation, recurring fixed and variable cost.

Expenses will be considered as an when they are occurred and will be part of cost for the project to the company.

After calculating the cost we should calculate the total revenue and deduct incurred cost and taxes, it will give us the net cash flow and discounting it with the cost of capital will give us Net present value of Cash Flow.

Below is the excel formula to calculate PV:

=PV (rate, nper, pmt, fv)

  • rate – Cost of capital.
  • nper - The total number of payment periods.
  • pmt - The payment made each period.
  • fv – Total cash flow

If NPV is positive the project is profitable and company may consider it.

Below is the calculation for present value of the future cash flows on the excel sheet.

Year

0

1

2

3

4

Per Unit Cost

10

10

15

10

Number   of Unites

5000

6000

2000

3000

Price per unit

25

25

15

40

Incremental Revenue

125000

150000

30000

120000

Incremental Variable cost /Expenses

50000

60000

30000

30000

Incremental Fixed cost /Expenses

2000

2000

2000

2000

Implementation Cost (fixed one time)

3500

0

0

0

0

Total Cost

3500

52000

62000

32000

32000

Total Cash Flow

-3500

73000

88000

-2000

88000

Tax

0

400

500

0

500

Net Cash Flow after Tax

-3500

72600

87500

-2000

87500

Cost of Capital

8%

8%

8%

8%

8%

PV of Cash Flow

₹ -3,500.00

₹ 67,222.22

₹ 75,017.15

₹ -1,587.66

₹ 64,315.11

NPV

₹ 2,01,466.82

Here is the worksheet with showing formula in the cell.

Capital Budgeting
Year 0 1 2 3 4
Per Unit Cost 10 10 15 10
Number   of Unites 5000 6000 2000 3000
Price per unit 25 25 15 40
Incremental Revenue =C5*C4 =D5*D4 =E5*E4 =F5*F4
Incremental Veriable cost /Expenses =C3*C4 =D3*D4 =E3*E4 =F3*F4
Incremental Fixed cost /Expenses 2000 2000 2000 2000
Implimentation Cost (fixed one time) 3500 0 0 0 0
Total Cost =B10+B11+B12 =C10+C11+C12 =D10+D11+D12 =E10+E11+E12 =F10+F11+F12
Total Cash Flow =B8-B13 =C8-C13 =D8-D13 =E8-E13 =F8-F13
Tax 0 400 500 0 500
Net Cash Flow after Tax =B16-B17 =C16-C17 =D16-D17 =E16-E17 =F16-F17
Cost of Capital 0.08 0.08 0.08 0.08 0.08
PV of Cash Flow =-PV(B19,B2,0,B18) =-PV(C19,C2,0,C18) =-PV(D19,D2,0,D18) =-PV(E19,E2,0,E18) =-PV(F19,F2,0,F18)
NPV =SUM(B21:F21)